A block of mass 6.69 kg is on a horizontal surface. The coefficient of kinetic friction between the block and the surface is 0.159.
a) What horizontal force would be required to keep the block moving at a constant velocity?
b) What horizontal force is required to accelerate the block at 4.67 m/s^2 ?
1
Expert's answer
2019-10-21T09:06:27-0400

The Newton's second law gives


"ma=F-F_{\rm frict}""ma=F-\mu mg"

(a) when a block miving with constant velocity its acceleration is zero , so


"F=\mu mg=0.159\times 6.69\:\rm kg\times 9.81\:\rm m/s^2=10.4\: N"

(b) if acceleration of the block is "4.67\: \rm m/s^2", we obtain


"F=ma+\mu mg""=6.69\:\rm kg\times4.67\: \rm m/s^2+10.4\: N =41.7\: N"